# MARS-Sep: Multimodal-Aligned Reinforced Sound Separation

### v1

@@ Appendix B Experimental Setup
Following (Dong et al., 2023; Cheng et al., 2025b), for all audio samples, we conducted experiments on samples of length 65535 (approximately 4 seconds) at a sampling rate of 16 kHz. For spectrum computation, we employed a short-time Fourier transform (STFT) with a filter length of 1024, a hop length of 256, and a window size of 1024. All images were resized to $224\times 224$ pixels. The audio model in this paper is a widely used 7-layer U-Net network with $k=32$, generating 32 intermediate masks. All models were trained with a batch size of 128, using the Adam optimizer with parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\sigma=10^{-8}$, for 200,000 steps. Additionally, we employed warm-up and gradient clipping strategies, following (Dong et al., 2023). We compute the signal-to-distortion ratio (SDR) using museval (Stöter et al., 2018). All experiments were conducted on a single A100 GPU with 40GB display memory.
@@ E.1 Ablation Studies
E.1.1 Effect of Progressive Fine-tuning on Source Discrimination of ImageBind
@@
E.1.2 Effect of reinforcement learning and encoder fine-tuning under different training pipelines.
@@ E.3 Efficacy of the Progressive Alignment Finetuning
However, for image query separation, the improvement is less consistent - the no-fine-tuning model performs competitively on SDR and SAR, suggesting that visual embeddings pre-trained on large-scale video data are already well-aligned with audio representations in VGGSOUND-clean+. Nonetheless, fine-tuning slightly boosts the CLAP similarity, implying better semantic alignment between modalities.
When all modalities are jointly composed as omni-modal queries, the three-stage fine-tuning strategy again achieves the best balance between separation quality and semantic consistency, achieving the highest SI-SDRi and $\text{CLAP}_{t}$ scores. Overall, these results confirm that while the benefits of progressive alignment may vary across modalities, it remains most effective for improving cross-domain generalization and semantic consistency in complex, in-the-wild sound mixtures.
E.4 Additional Qualitative Results in the TQSS Setting

### v2

@@ Appendix B Experimental Setup
Following (Dong et al., 2023; Cheng et al., 2025d), for all audio samples, we conducted experiments on samples of length 65535 (approximately 4 seconds) at a sampling rate of 16 kHz. For spectrum computation, we employed a short-time Fourier transform (STFT) with a filter length of 1024, a hop length of 256, and a window size of 1024. All images were resized to $224\times 224$ pixels. The audio model in this paper is a widely used 7-layer U-Net network with $k=32$, generating 32 intermediate masks. All models were trained with a batch size of 128, using the Adam optimizer with parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\sigma=10^{-8}$, for 200,000 steps. Additionally, we employed warm-up and gradient clipping strategies, following (Dong et al., 2023). We compute the signal-to-distortion ratio (SDR) using museval (Stöter et al., 2018). All experiments were conducted on a single A100 GPU with 40GB display memory.
@@ E.1 Ablation Studies
This section provides a systematic analysis of the key factors influencing MARS-Sep, including hyperparameters, modality fusion strategies, and encoder characteristics. Across all experiments, we observe a consistent pattern: stable policy optimization and semantically aligned reward modeling jointly determine final separation quality, and different components contribute in a complementary way.
E.1.1 Hyperparameter Settings
Tables 4-7 summarize the effect of several hyperparameters in omni-queried cases. The concentration parameter $\kappa$ of the Beta policy introduces a direct trade-off between exploration and stability: small values yield overly noisy samples, while large values cause premature policy collapse. $\kappa=9$ achieves the best balance between semantic reward optimization and signal fidelity. The PPO clip range $\epsilon$ exhibits very limited sensitivity, indicating that the trust-region update is already sufficiently stable.
In contrast, the entropy coefficient $\lambda_{H}$ and the KL coefficient $\lambda_{KL}$ noticeably influence exploration behavior. Moderate entropy regularization encourages continued mask diversity, while a lightweight KL penalty prevents the policy from drifting too far from its previous snapshot, which stabilizes reward estimation. Overall, these trends align with observations from GRPO-style RL training in large models—mild exploration and gentle update constraints are most effective for maintaining semantic consistency.
$\kappa$ 3 6 9 (Default) 12 15
Mean SDR 6.99±0.71 7.84±0.74 7.93±0.75 7.74±0.75 7.70±0.77
Mean $\textbf{CLAP}_{t}$ 8.41±0.90 8.75±0.92 9.22±0.90 9.19±0.94 9.10±0.88
Table 4: Effect of Beta distribution concentration parameter ($\kappa$).
$\epsilon$ 0.01 0.1 0.2 (Default) 0.3 0.4
Mean SDR 7.91±0.75 7.90±0.74 7.93±0.75 7.91±0.75 7.91±0.73
Mean $\textbf{CLAP}_{t}$ 9.00±0.95 9.08±0.92 9.22±0.90 9.14±0.88 9.12±0.88
Table 5: Effect of PPO clip range parameter ($\epsilon$).
$\lambda_{H}$ 0.01 0.1 0.2 (Default) 0.3 0.4
Mean SDR 7.20±0.74 7.68±0.72 7.93±0.75 7.62±0.72 7.57±0.77
Mean $\textbf{CLAP}_{t}$ 8.50±0.89 8.76±0.93 9.22±0.90 8.79±0.92 8.46±1.01
Table 6: Effect of Entropy coefficient parameter ($\lambda_{H}$).
$\lambda_{KL}$ 0 0.1 (Default) 0.2
Mean SDR 7.43±0.72 7.93±0.75 7.69±0.70
Mean $\textbf{CLAP}_{t}$ 9.63±0.90 9.22±0.90 9.19±0.92
Table 7: Effect of KL Divergence coefficient parameter ($\lambda_{KL}$).
E.1.2 Modality Aggregation Module
We compared several query fusion mechanisms beyond MLBP, including Max Pooling, Average Pooling, and Learnable Weighted Sums. Results are displayed in Table 8 and Table 9. Although Learnable Weighted Sums slightly outperform MLBP in the Audio-Query and Omni-Query settings of VGGSOUND-clean+, MLBP exhibits more stable behavior in Text-Query and Image-Query scenarios and achieves the strongest and most consistent gains on the cross-domain MUSIC-clean+ benchmark.
This advantage arises because MLBP explicitly models multiplicative interactions across modalities—especially between text and vision—allowing the reward model to form a more reliable cross-modal semantic anchor. As a result, MLBP proves particularly effective when semantic cues are complex or span multiple modalities, outperforming simpler pooling or weighting schemes under these challenging conditions.
Table 8: Comparison of sound separation performance among different methods on VGGSOUND-clean+ dataset.
Methods VGGSOUND-clean+
[…]
Text Query Sound Separation
Max Pooling 6.83±0.68 8.69±1.00 14.19±0.75 4.35±0.44 8.99±0.92
Learned Weighted Sums 6.79±0.65 8.95±1.00 13.77±0.75 4.34±0.46 8.95±0.95
Average Pooling 6.75±0.66 9.20±1.00 13.54±0.75 9.18±0.79 8.97±0.94
MLBP (ours) 6.91±0.68 9.14±1.00 13.73±0.77 4.55±0.44 9.03±0.94
Audio Query Sound Separation
Max Pooling 7.11±0.61 11.21±0.99 11.72±0.79 4.39±0.53 8.67±0.93
Learned Weighted Sums 7.23±0.65 11.57±0.99 11.82±0.82 4.24±0.55 8.75±0.97
Average Pooling 6.69±0.61 11.76±1.02 11.12±0.76 3.90±0.60 8.42±0.91
MLBP (ours) 7.33±0.67 11.63±1.00 12.00±0.84 4.36±0.50 8.91±0.91
Image Query Sound Separation
Max Pooling 6.78±0.66 9.52±1.04 13.76±0.74 4.43±0.49 8.68±0.90
Learned Weighted Sums 7.15±0.69 10.03±1.07 13.95±0.75 4.80±0.50 9.04±0.92
Average Pooling 7.03±0.68 10.20±1.07 13.95±0.76 4.69±0.48 8.81±0.91
MLBP (ours) 6.93±0.67 10.18±1.04 13.41±0.72 4.57±0.47 9.19±0.91
Composed Omni-modal Query Sound Separation
Max Pooling 7.86±0.72 10.31±1.00 14.54±0.92 5.13±0.45 8.94±0.92
Learned Weighted Sums 8.08±0.75 10.58±1.00 14.74±0.95 5.32±0.47 9.07±0.95
Average Pooling 7.87±0.74 10.96±1.03 14.52±0.91 5.20±0.46 9.18±0.94
MLBP (ours) 7.93±0.75 10.65±1.00 14.49±0.95 5.20±0.45 9.22±0.90
Methods MUSIC-clean+
[…]
Text Query Sound Separation
Max Pooling 12.80±0.92 17.37±1.20 18.58±0.89 9.55±0.81 5.34±0.89
Learned Weighted Sums 12.60±0.82 17.79±1.21 18.23±0.84 9.47±0.80 5.39±0.92
Average Pooling 12.08±0.89 17.31±1.19 17.73±0.87 8.82±0.85 5.14±0.92
MLBP (ours) 12.91±0.93 17.61±1.17 18.28±0.93 9.85±0.82 6.18±0.93
Audio Query Sound Separation
Max Pooling 11.13±0.87 18.72±1.14 15.56±0.84 7.81±1.06 5.38±0.96
Learned Weighted Sums 10.98±0.89 18.92±1.10 14.69±0.90 8.03±1.00 5.77±0.98
Average Pooling 10.48±0.90 18.69±1.03 14.16±0.90 7.62±1.08 5.46±1.00
MLBP (ours) 11.73±0.88 19.65±1.14 15.25±0.86 8.38±1.03 5.64±1.06
Image Query Sound Separation
Max Pooling 13.29±1.05 18.69±1.20 18.23±1.02 10.35±0.95 6.44±1.05
Learned Weighted Sums 13.21±0.93 18.75±1.17 18.23±0.95 10.31±0.92 6.09±1.01
Average Pooling 12.80±0.96 18.27±1.18 17.85±0.97 9.82±0.93 6.36±1.03
MLBP (ours) 13.64±1.06 19.24±1.16 18.05±1.06 10.70±0.89 6.94±1.06
Composed Omni-modal Query Sound Separation
Max Pooling 13.61±1.01 19.61±1.21 18.10±0.97 10.58±0.88 6.61±0.96
Learned Weighted Sums 13.41±0.90 19.54±1.18 17.99±0.90 10.43±0.81 6.49±0.95
Average Pooling 13.15±0.95 19.21±1.23 17.73±0.92 10.09±0.84 6.46±0.94
MLBP (ours) 13.89±0.98 19.90±1.18 17.99±0.97 10.78±0.81 6.82±0.99
E.1.3 Effect of Progressive Fine-tuning on Source Discrimination of ImageBind
@@
E.1.4 Effect of reinforcement learning and encoder fine-tuning under different training pipelines.
@@ E.3 Efficacy of the Progressive Alignment Finetuning
While the improvement from progressive fine-tuning is slightly less pronounced in the image-query setting, this can be naturally explained by the fact that ImageBind’s visual encoder is pretrained on large-scale visual data and already provides a stable embedding space. Thus, image-audio alignment requires less additional adaptation than text-audio or audio-audio alignment. In contrast, the text and audio branches differ more significantly in semantic granularity and distribution, making them more sensitive to staged alignment.
Nonetheless, the three-stage strategy consistently provides the most robust performance across text, audio, and omni-modal queries, where semantic compositionality is more complex. These results confirm that progressive alignment is especially beneficial for cross-modal generalization and semantic consistency in challenging, in-the-wild mixtures.
E.4 Memory Consumption and Inference Latency
To validate the training time and memory consumption and inference efficiency, we conduct additional experiments to measure training compute, hardware configuration, memory usage, and inference latency on standard hardware. Table 14 summarizes the results. Both OmniSep and our MARS-Sep model were trained on an Nvidia A800 40GB GPU, and the memory consumption for all modalities remains comparable at approximately 35.5 GB. In terms of training efficiency, MARS-Sep requires roughly 8 hours per epoch with a batch size of 128 and 10k steps, which is about 50% of the baseline efficiency reported by OmniSep (about 4 hours per epoch). Despite the increased training cost, our model introduces virtually no additional inference overhead: the real-time factors (RTF) across text, image, audio, and omni-modality inputs remain on par with OmniSep, with differences within normal variance. These results demonstrate that MARS-Sep maintains comparable inference-time efficiency while achieving its improvements through a modest increase in training-time computation.
Method Hardware Memory Consumption Training Time per Epoch (Batch Size=128, 10k steps) Inference Latency per Batch(RTF) (Batch Size = 4)
OmniSep $\sim$ 35.5GB Nvidia A800 40GB $\sim$ 4 hours Text: 0.1283s Image: 0.0841s Audio: 0.0812s Omni: 0.0801s
MARS-Sep (Ours) $\sim$ 35.5GB Nvidia A800 40GB $\sim$ 8 hours Text: 0.1202s Image: 0.0816s Audio: 0.0814s Omni: 0.0893s
Table 14: Efficiency reporting on standard hardware. Inference Latency is tested on VGGSOUND dataset.
E.5 User Study for Evaluating Cross-Modal Semantic Alignment
To complement the CLAP-based automatic metrics and address concerns about potential biases inherited from CLAP, we conducted a human perceptual study to directly evaluate the semantic alignment between user queries and the separated audio. We sampled 100 query-audio pairs from the test set, covering diverse sound events and including both text, image and audio queries. For each pair, we presented participants with the query and two audio excerpts-one target audio and one separated by our method or OmniSep-in randomized order.
10 non-expert listeners participated in the study, each providing 100 judgments. Participants rated the semantic consistency between the query and each audio sample on a 1-5 Likert scale, and additionally chose the sample that better matched the query in a pairwise preference setting. As shown in Table 15, MARS-Sep achieves higher semantic-matching scores and is preferred against the target audio more often than OmniSep across all three query conditions. These results indicate that our improvements are not merely artifacts of CLAP similarities but are also perceived by human listeners, supporting the claim that MARS-Sep produces more semantically aligned separations.
Method Semantic Match (1-5) Pairwise Preference (%)
Target Audio 4.76 $\pm$ 0.20 /
OmniSep 3.42 $\pm$ 0.51 15.4
MARS-Sep (Ours) 3.57 $\pm$ 0.48 23.2
Table 15: Human user study evaluating semantic alignment between queries and separated audio.
E.6 Additional Qualitative Results in the TQSS Setting